oint Cost Allocation—Weighted Average Method

Gary’s Grooves Co. produces two types of carving knives, one with a handle made of a polymer that looks like walnut wood and another with a handle made with a polymer that looks like red oak. The knives are made through a joint production molding process that produces 344 knife blades for red oak handle knives and 172 knife blades for walnut handle knives at the split-off point. The polymer for the red oak handle knife blades requires twice as much cooling time as the polymer for the walnut handle knife blades, although all knives are removed from the joint molding process at the same time (i.e., once the cooling for the red oak handle knives is complete). The joint production process costs a total of $5,000.

Assuming the company allocates joint costs using the weighted average method based on the required cooling time of the two joint products, determine the amount of joint production costs allocated to each type of knife using the weighted average method.

Joint Product Allocation
Red oak handle carving knife $
Walnut handle carving knife
Totals $
Answer #1

Weighted production units:

Red Oak = 344 X 2 = 688

Walnut = 172 X 1 = 172

Total = 688 + 172 = 860

Joint cost allocation:

Red Oak = $5,000 X 688/860 = $4,000

Walnut = $5,000 X 172/860 = $1,000

Joint product Allocation
Red oak handle carving knife $4,000
Walnut handle carving knife $1,000
Total $5,000